OpenStudy (anonymous):

Simplify sin(arccos(x/(x^2) + 1))

OpenStudy (anonymous):

aye try mathway.com/ they should have the answer on there

OpenStudy (anonymous):

\[\sin\left[\cos^{-1}\left(\frac{x}{x^2+1}\right)\right]~~?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay. Suppose \(\theta=\cos^{-1}\left(\dfrac{x}{x^2+1}\right)\). Then \(\cos\theta=\dfrac{x}{x^2+1}\), right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Consider a triangle that agrees with this:  Makes sense so far?